1. A container is filled with 16.0 g of O2 and 14.0 g of N2.

a. What is the volume of the container at STP?

b. What is the partial pressure of the O2 gas?

c. What is the mole fraction of N2 in the mixture?

2. Look carefully at the following reaction: N2(g) + 3H2(g) -> 2NH3(g)

How will the enthalpy change (H) compared to the internal energy change (E) for this reaction? That is, will H be equal to, greater than, or less than E? Explain.
Answer #1

mol of O2 = mass/MW = 16/32 = 0.5

mol of N2 = mass/MW = 14/28 = 0.5

Total mol = 0.5+0.5 = 1

At STP. 1 mol occupies 22.4 L of space

therefore

a) STP ratio = 22.4L

b)

P of O2

Px = xO2*PT

Px = 0.5*1atm = 0.5 atm of O2

c)

mole fraction of N2 = mol of N2 / total mol = 0.5/1 = 0.5

2)

N2(g) + 3H2(g) -> 2NH3(g)

H = U + PV

dH = dU + dPdV

from this example, since number of gas mole decrease drastically (from 4 to 2), then expect dV (to be negative)

Then dH = dU - dP*dV

E reaction decreases